# Spike-frequency and h-current based adaptation are dynamically equivalent in a Wilson-Cowan field model

### v2

@@ Acknowledgements.
We thank Prof. Tilo Schwalger (Technische Universität Berlin) for very helpful comments and discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG) as part of the CRC 1315 (project number 327654276).
@@ A.1 Analysis of the Wilson-Cowan model for zero and small values of $b$
For $\vec{X}=\begin{bmatrix}\hat{\Phi}\\ \hat{\varphi}\end{bmatrix}$, Equations (14) can be written in vector notation,
@@
$\begin{split}A_{2}(k)=&\begin{bmatrix}\frac{1}{\tau_{e}}(-1+w_{ee}F^{\prime}_{e}(B_{e})\hat{w}_{e}(k))&-\frac{w_{ei}}{\tau_{e}}F^{\prime}_{e}(B_{e})\hat{w}_{i}(k)\\ \frac{w_{ie}}{\tau_{i}}F^{\prime}_{i}(B_{i})\hat{w}_{e}(k)&\frac{1}{\tau_{i}}(-1-w_{ii}F^{\prime}_{i}(B_{i})\hat{w}_{i}(k))\end{bmatrix}\\ B_{l}=&\ w_{le}\tilde{u}_{e}-w_{li}\tilde{u}_{i}+I_{l},\ l\in\{e,i\},\end{split}$ (16)
@@
$\begin{split}\det(A_{2}(k_{0}))>0\\ \text{tr}(A_{2}(k_{0}))<0\end{split}$ (17)
@@ A.2 Static and dynamic Turing bifurcations for small values of $b$
Without adaptation, the system shows static Turing bifurcations. For a small value of b=0.2b=-0.2, nearly all formerly static Turing bifurcations are either dynamic, or both dynamic and static Turing bifurcations occur at different non-zero wavenumbers as, for example, in Figure 4, upper row, for $b\neq 0$, indicated by small arrows in panels below plots.
@@ A.3 Activity traces for traveling and standing waves with finite adaptation
Figure 8: Activity patterns corresponding to the Turing-unstable down-state shown in Figure 4, bottom row, for $x\in[-50,50]$.
Figure 8 show the excitatory activity for very weak adaptation. While for b=0.03b=-0.03 the waves travel over space, they becomes stationary for b=0.02b=-0.02
@@ B.3 Regularity of patterns
We characterized the spatio-temporal pattern using the Kuramoto order parameter. We computed the analytical signal representation of the activity $u_{e}(x,t)$ using the Hilbert-Transform
@@
we calculate the Kuramoto order parameter $R(t)$ for each time step $t$
$R(t)=\left|\frac{1}{n}\sum_{x}e^{i\theta(x,t)}\right|$ (20)
@@
$r=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(R(t)-\bar{R})}\ ,\ \ \text{ }\ \bar{R}=\frac{1}{T}\sum_{t=1}^{T}R(t)$ (21)
over a time interval $[0,T]$. We considered patterns to be regular if $r\leq 10^{-2}$. This threshold corresponds to the mean over the ten most irregular patterns found when compiling Figure 5A.
@@ B.4 Phase velocity and speed
To compute the speed in $\bigl[\frac{unit_{x}}{unit_{t}}\bigr]$, where $unit_{x}=\frac{1}{dx}$, of a spatio-temporal pattern, we used the approach of [78]. We first determine the analytical representation of the activity $u_{e}(x,t)$ as in Equation (18), where $\theta(x,t)$ is the instantaneous phase at position $x$ and time step $t$ as in Equation (19). The local speed $s(x,t)$ is computed from the temporal and spatial derivatives $\frac{\partial\theta}{\partial x}(x,t)$ of the instantaneous phase,
@@
Since spatio-temporal patterns may be composed of traveling waves of different speeds (see, e.g. see Figure 6B, panel 5), we apply a similar approach as for the dominant spatial frequency. We compute a histogram over all local speed values appearing per activity pattern and take the speed value appearing the most often. In Figure 9, we illustrate this procedure for four different activity patterns.
B.5 Visualization of feature-spaces with initialization close to zero
Figure 10: Properties of the spatio-temporal patterns emerging in Hopf- and Turing-unstable states for increasing h-current strength. A Wavenumber $k_{0}$ (top row), dominant spatial frequency $f_{x}$ (middle row), and dominant temporal frequency $f_{t}$ (bottom row) for each location in state space. The figure shows slices of state space spanned by the external input $(I_{e},I_{e})\in[-2.5,2]\times[-3.5,1]$ for $b\in\{0,-0.5,-1.0\}$. Purple colors denote the corresponding feature values. Darker colors indicate higher values, see color bar. Yellow denotes states of no spatio-temporal activity, light green denotes states with irregular patterns with a standard deviation $r$ of the Kuramoto-order parameter of $r>10^{-2}$. Thin lines denote the boundaries of the dynamical regimes shown in Figure 2A. B Histograms and kernel density estimates (KDE; thin lines) of the numerically calculated spatial frequency ($f_{x}$; top row), the spatial frequency corresponding to $k_{0}$ ($\frac{L\cdot k_{0}}{2\pi}$; middle row), and the spatial frequency corresponding to $k_{max}$ ($\frac{L\cdot k_{max}}{2\pi}$; bottom row) in the regimes of Hopf instability (left column), Turing instability in down-states (middle column), and Turing instability in up-states (right column) for b=0.5b=-0.5 Dashed vertical lines with numbers denote the location of the peak of the KDE. Numerical simulations were conducted with initializations close to zero.
B.6 Input pair tables for Figures 2 and 6
Table 3: Input pairs $(I_{e},I_{i})$ for Figure 2.
@@ B.6 Input pair tables for Figures 2 and 6
Table 4: Input pairs $(I_{e},I_{i})$ for Figure 6.
@@ B.7 Spatial frequencies of the spatio-temporal patterns emerging for different initializations
Figure 11: A Activity traces 4, 6, and 3 (top to bottom row) of Figure 2 initialized close to the fixed points of Turing instability in a down-state (first row), in an up-state (second row), and of Hopf instability but with different seeds. B Consistency of the spatial dynamics for 30 different seeds for each activity in A. Visualized are interquartile ranges (boxes), medians (horizontal lines inside boxes), spread (error bars), and outliers outside of spread (circles) for the Turing-unstable down- (green), and up-state (purple), and the Hopf-unstable state (red).

### v3

@@ Acknowledgements.
We thank Prof. Tilo Schwalger (Technische Universität Berlin) for very helpful comments and discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG) as part of the CRC 1315 (project number 327654276). It was published on arXiv as a preprint in Oct. 2025 [77].
@@ A.1 Analysis of the Wilson-Cowan model for zero and small values of $b$
For $\vec{X}=\begin{bmatrix}\hat{\Phi}\\ \hat{\varphi}\end{bmatrix}$, Eqs. (10) can be written in vector notation,
@@
$\begin{split}A_{2}(k)=&\begin{bmatrix}\frac{1}{\tau_{e}}(-1+w_{ee}F^{\prime}_{e}(B_{e})\hat{w}_{e}(k))&-\frac{w_{ei}}{\tau_{e}}F^{\prime}_{e}(B_{e})\hat{w}_{i}(k)\\ \frac{w_{ie}}{\tau_{i}}F^{\prime}_{i}(B_{i})\hat{w}_{e}(k)&\frac{1}{\tau_{i}}(-1-w_{ii}F^{\prime}_{i}(B_{i})\hat{w}_{i}(k))\end{bmatrix}\\ B_{l}=&\ w_{le}\tilde{u}_{e}-w_{li}\tilde{u}_{i}+I_{l},\ l\in\{e,i\},\end{split}$
@@
$\begin{split}\det(A_{2}(k_{0}))>0\\ \text{tr}(A_{2}(k_{0}))<0\end{split}$
@@ A.2 Static and dynamic Turing bifurcations for small values of $b$
Figure 7 shows the state spaces for $b=0$ and b=0.2b=-0.2 including the distinction of static vs. dynamic Turing bifurcations. Without adaptation, the system shows static Turing bifurcations. For a small value of b=0.2b=-0.2, nearly all formerly static Turing bifurcations are either dynamic, or both dynamic and static Turing bifurcations occur at different non-zero wavenumbers as, for example, in Fig. 4, upper row, for $b\neq 0$, indicated by small arrows in panels below plots.
@@ A.3 Activity traces for traveling and standing waves with finite adaptation
Figure 8: Activity traces $u_{e}(x,t)$ corresponding to the Turing-unstable down-state shown in Fig. 4, bottom row, for $x\in[-50,50]$.
Figure 8 show the excitatory activity for very weak adaptation. While for b=0.03b=-0.03 the waves travel over space, they become stationary for b=0.02b=-0.02
@@ B.3 Regularity of patterns
We characterized the spatio-temporal pattern using the Kuramoto-order parameter. We computed the analytical signal representation of the activity $u_{e}(x,t)$ using the Hilbert-Transform
@@
we calculate the Kuramoto-order parameter $R(t)$ for each time step $t$
$R(t)=\left|\frac{1}{n}\sum_{x}e^{i\theta(x,t)}\right|$
@@
$r=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(R(t)-\bar{R})}\ ,\ \ \text{ }\ \bar{R}=\frac{1}{T}\sum_{t=1}^{T}R(t)$
over a time interval $[0,T]$. We considered patterns to be regular if $r\leq 10^{-2}$. This threshold corresponds to the mean over the ten most irregular patterns found when compiling Fig. 5A.
@@ B.4 Phase velocity and speed
To compute the speed in $\bigl[\frac{unit_{x}}{unit_{t}}\bigr]$, where $unit_{x}=\frac{1}{dx}$, of a spatio-temporal pattern, we used the approach of [79]. We first determine the analytical representation of the activity $u_{e}(x,t)$ as in Eq. (12), where $\theta(x,t)$ is the instantaneous phase at position $x$ and time step $t$ as in Eq. (13). The local speed $s(x,t)$ is computed from the temporal and spatial derivatives $\frac{\partial\theta}{\partial x}(x,t)$ of the instantaneous phase,
@@
Since spatio-temporal patterns may be composed of traveling waves of different speeds (e.g., see Fig. 6B, panel 5), we apply a similar approach as for the dominant spatial frequency. We compute a histogram over all local speed values appearing per activity pattern and take the speed value appearing the most often. In Fig. 9, we illustrate this procedure for four different activity patterns.
B.5 Input pair tables for Figures 2 and 6
Table 3: Input pairs $(I_{e},I_{i})$ for Fig. 2.
@@ B.5 Input pair tables for Figures 2 and 6
Table 4: Input pairs $(I_{e},I_{i})$ for Fig. 6.
@@
B.6 Visualization of feature-spaces with initialization close to zero
Figure 10: Properties of the spatio-temporal patterns emerging in Hopf- and Turing-unstable states for increasing h-current strength. A Wavenumber $k_{0}$ (top row), dominant spatial frequency $f_{x}$ (middle row), and dominant temporal frequency $f_{t}$ (bottom row) for each location in state space. The figure shows slices of state space spanned by the external input $(I_{e},I_{e})\in[-2.5,2]\times[-3.5,1]$ for $b\in\{0,-0.5,-1.0\}$. Purple colors denote the corresponding feature values. Darker colors indicate higher values, see color bar. Yellow denotes states of no spatio-temporal activity, light green denotes states with irregular patterns with a standard deviation $r$ of the Kuramoto-order parameter of $r>10^{-2}$. Thin lines denote the boundaries of the dynamical regimes shown in Fig. 2A. B Histograms and kernel density estimates (KDE; thin lines) of the numerically calculated spatial frequency ($f_{x}$; top row), the spatial frequency corresponding to $k_{0}$ ($\frac{L\cdot k_{0}}{2\pi}$; middle row), and the spatial frequency corresponding to $k_{max}$ ($\frac{L\cdot k_{max}}{2\pi}$; bottom row) in the regimes of Hopf instability (left column), Turing instability in down-states (middle column), and Turing instability in up-states (right column) for b=0.5b=-0.5 Dashed vertical lines with numbers denote the location of the peak of the KDE. Numerical simulations were conducted with initializations close to zero.
@@ B.7 Spatial frequencies of the spatio-temporal patterns emerging for different initializations
Figure 11: A Activity traces 4, 6, and 3 (top to bottom row) of Fig. 2 initialized close to the fixed points of Turing instability in a down-state (first row), in an up-state (second row), and of Hopf instability but with different seeds. B Consistency of the spatial dynamics for 30 different seeds for each activity in A. Visualized are interquartile ranges (boxes), medians (horizontal lines inside boxes), spread (error bars), and outliers outside of spread (circles) for the Turing-unstable down- (green) and up-state (purple), and the Hopf-unstable state (red).